OpenStudy (hockeychick23):

Suppose a binomial event has a probability of success of 0.1 and 750 trials are performed. What is the standard deviation of the possible outcomes?

OpenStudy (anonymous):

st.dev.=sqrt( n*p*(1-p) ) So you have p=0.1 and n=750 st.dev=sqrt( 750* 0.1*0.9)

OpenStudy (dan815):

uhh here is a more natural approach if u dont wanna believe in the formula yet

OpenStudy (dan815):

you know that in 1 standard deviation about 68% of the outcomes have to exist closed to the mean

OpenStudy (hockeychick23):

oh wow that was actually a lot quicker than i thought, thanks, and yep

OpenStudy (dan815):